# Unified Theory of Magnetization Temperature Dependence in Ferrimagnetics

### v1

@@ Appendix A Primary reasons for low magnon damping in YIG
•  Firstly, there is no magnon-phonon coupling. Fe3+ ions have a half-filled orbital, which means that the orbital angular momentum $L=0$. As a result, the spin-orbit interaction, which is proportional to $\bm{S}\cdot\bm{L}$ and is responsible for magnon-phonon coupling, is also suppressed. Consequently, the magnon damping, which occurs due to their decay into phonons and magnons, is also eliminated.
@@
Appendix B Positions of Fe3+ ions in elementary cell of YIG
To define the number of nearest neighbors and the relevant distances to assign the coordination spheres, we have drawn in Fig. 1 (b) the Fe3+ ions, oriented such that (a)-Fe3+ ions are placed in the corners and the center of the half-cube. The particular Fe3+ ion located at (a)-position (0,0,0) (see coordinate labels in Fig. 1 (a)) has $Z_{\rm aa}=8$ nearest-neighbor (a)-Fe3+ ions in positions (±$(\pm\frac{1}{4},\pm\frac{1}{4},\pm\frac{1}{4})$, separated by distance $\Delta_{\rm aa}=\dfrac{\sqrt{3}}{4}\approx 0.433$. The coordinates of the nearest $Z_{\rm ad}=6$ neighboring sites (d) with respect to the (a)-Fe3+ ion at $(0,0,0)$ are ($0,\frac{1}{4},-\frac{1}{8}$),  ($0,-\frac{1}{4},\frac{1}{8}$);  ($-\frac{1}{8},0,-\frac{1}{4}$),  ($\frac{1}{8},0,\frac{1}{4}$); ($\frac{1}{4}$, $-\frac{1}{8}$, 0), ($-\frac{1}{4},\frac{1}{8},0$), with the separation $\Delta_{\rm ad}=\dfrac{\sqrt{5}}{8}\approx 0.2795$.
To find the environment of the (d)-Fe3+ ion, we place it at the origin. Then, there are $Z_{\rm dd}=4$ nearest (d)-neighbors with the coordinates ($-\frac{1}{8},\frac{1}{8},-\frac{1}{4}$), ($-\frac{1}{4},\frac{1}{8},-\frac{1}{8}$),($-\frac{1}{4},-\frac{1}{8},\frac{1}{8}$) and ($\frac{1}{8},\frac{1}{8},\frac{1}{4}$) at the distance $\Delta_{\rm dd}=\dfrac{\sqrt{6}}{8}\approx 0.306$.  This central (d)-Fe3+ ion has $Z_{\rm da}=4$ (a)-Fe3+ neighbors at positions ($0,-\frac{1}{8},-\frac{1}{4}$),(0,$-\frac{1}{8},\frac{1}{4}$),($\frac{1}{4},\frac{1}{8},0$), (-$\frac{1}{4},\frac{1}{8},0$) with the separation $\Delta_{\rm da}=\Delta_{\rm ad}=\dfrac{\sqrt{5}}{8}\approx 0.2795$.
Appendix C Exchange integrals
@@
6 Princep [25] 2017 Improved Neutron scattering of SW -79    0.0 |||| -12.8 -6.0 1018 -0.82 86
@@
Table 1: Imporant parameters of YIG. The columns from left to right: the reference to the data origin and the method used to extract it; the exchange integrals $J_{ij}$ (values in K), the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF}}$ in the mean-field approximation computed using Eq. (16a); the ratio $\overline{S}_{\rm d}^{z}/\overline{S}_{\rm a}^{z}$ for $T\to T_{{}_{\rm C}}^{{}^{\rm MF}}$ from below, Eq. (16b); the exchange frequency $\omega_{\rm ex}$ (in K) according to (34c). Estimates of $\omega_{\rm ex}$ obtained by various authors, using a number of techniques, vary from 40 to 100 K, see Tab. 1 in [50]. The experimental value of YIG Curie temperature is $T_{{}_{\rm C}}^{\rm Exp}\approx 559\,$K. For the comparison with other published data, note that 111\,1meV $\approx 11.6\,$K.
@@
Appendix D Approximations for magnon frequencies $\omega_{j}(k)$
@@
$\begin{split}\omega(k)=&\,\,\omega_{\rm ex}(ak)^{2}\,,\\ \omega_{\rm ex}=\,&\frac{5}{16}(8J_{\rm aa}+3J_{\rm dd}-5J_{\rm ad})\approx 80\,\mbox{K}\,,\end{split}$ (34c)
in agreement with the well known result [29]. However for $q>q_{*}$, above some relatively small transition value $q_{*}=1/\sqrt{25.4}\approx 0.2$, the dependence $\omega(k)$ becomes almost linear
$\omega_{\rm da1}\approx(-\omega_{\rm neg}+\omega_{\rm lin}\,ak)\,,$ (34d)
as seen in Fig. 3. In (34d), $\omega_{\rm neg}\approx 110\,$K and $\omega_{\rm lin}\approx 170\,$K. The nearly linear part of the dispersion laws $\omega_{\rm da1}$ and $\omega_{\rm ad1}$ originates from the fact that YIG is close to an antiferromagnetic arrangement  [^2] In YIG, the ratio of the numbers of spins oppositely oriented is 12/8= 3/2. Their difference 12-8 = 4 is small with respect to their half of the sum =10. 4/10 is quite a small number. In this sense, YIG magnetic structure is relatively close to AFM, where the dispersion laws exhibit a similar form [53].
@@
Appendix E Measuring Procedure and Apparatus
The magnetization measurements for this work were carried out using a Physical Property Measurement System (PPMS, Quantum Design), which allows temperature control from 400 K down to 1.8 K and enables the application of external magnetic fields of up to 9 T, generated by a pair of superconducting coils in Helmholtz configuration. The measurements were performed using the vibrating sample magnetometer (VSM) option integrated into the PPMS as an instrument insert.
A commercially obtained YIG sphere (Micro Spheres) with a diameter of 1 mm was fixed to the sample holder using nonmagnetic adhesive, with the easy axis—the $\langle 111\rangle$ crystallographic direction—aligned parallel to the applied external magnetic field. During the measurements, a constant external field of 1 T was applied, which is sufficient to ensure magnetic saturation of the YIG sphere across the entire available temperature range.
@@ Appendix E Measuring Procedure and Apparatus
Appendix F Numerical Methods
@@
For the fit shown in Fig. 2, represented by the violet curve ④, we used the experimental data from Ref. [21] [orange dots ②]. The high-temperature fit was based on the phenomenological expression $\overline{S}^{z}=A\cdot(T_{\rm C}-T)^{\beta}$, where $A$ and $\beta$ are fit parameters. The fitting was performed using the NonlinearModelFit function in Mathematica. To ensure a reliable result, it was necessary to carefully select the temperature range for the fit. Data in the interval from 350 K to 535 K was chosen. The upper bound had to remain below the Curie temperature $T_{\rm C}=560\,\text{K}$, as data too close to $T_{\rm{}_{C}}$ is affected by an experimental artifact—residual magnetization above $T_{\rm{}_{C}}$ caused by measurements taken in a relatively strong external magnetic field of 1 T. Conversely, the lower bound was set near room temperature, as the fitting model does not account for the saturation of magnetization at lower temperatures; thus, including such data would distort the fit parameters. It is worth noting that the fit results remained consistent under variations of the chosen interval limits by ±$\pm 10\,\text{K}$.
The curves $[-\overline{S}^{z}_{\mathrm{a}}(T)/\overline{S}^{z}_{\mathrm{a}}(0)]$ ⑨ and $\overline{S}^{z}_{\mathrm{d}}(T)/\overline{S}^{z}_{\mathrm{d}}(0)$ in Fig. 4, as well as their difference $\overline{S}^{z}(T)/\overline{S}^{z}(0)$ ⑤ in Figs. 2 and 4, were obtained by numerically solving the system of Eqs. (15) using the parameters specified in Eq. (14). The numerical solution was carried out using the FindRoot function in Mathematica.
For the curves ⑦ in Fig. 2 and the additional curves ⑪ and ⑫ in Fig. 5, Eq. (22a) was evaluated. This required calculating the total number of magnons within the first Brillouin zone as given by Eq. (22b), which involves an integration over the wavevector $\bm{k}$. In this work, we approximate the first Brillouin zone by a sphere of radius equal to the Debye wavevector $k_{\rm D}$, which for a BCC lattice is given by $k_{\rm D}=\sqrt[3]{12\pi^{2}/a^{3}}$, a simplification that provides a good approximation for our purposes. Accordingly, the integral defined in Eq. (23) over $n_{j}(\bm{k})$, Eq. (24), was performed in spherical coordinates, using the magnon dispersion relations $\omega_{j}(\bm{k})$ provided in Appendix D. The integration range was set from $k_{\rm min}=0$ to $k_{\rm max}=k_{\rm D}$. The numerical evaluation was carried out using the NIntegrate function in Mathematica.

### v2

@@ Appendix A Primary reasons for low magnon damping in YIG
• Firstly, there is no magnon-phonon coupling. Fe3+ ions have a half-filled orbital, which means that the orbital angular momentum $L=0$. As a result, the spin-orbit interaction, which is proportional to $\bm{S}\cdot\bm{L}$ and is responsible for magnon-phonon coupling, is also suppressed. Consequently, the magnon damping, which occurs as a result of their decay into phonons and magnons, is also eliminated.
@@
Appendix B Comparison of theoretical approaches to $M(T)$ dependence in single- and multi-sublattice magnetics
In this Appendix, we clarify the similarities and differences between our approach to EuO and EuS ferromagnetics, as presented in our previous paper [6], and our discussion of YIG developed in the current work.
EuO and EuS are the simplest objects available for theoretical description: they are ferromagnetic materials with only one magnetic ion in the elementary cell and have the smallest applicability parameter $1/Z$ for the mean field approximation (MFA) with $Z\approx 12$. These two aspects allowed us to develop a consistent step-by-step analysis of the temperature dependence of spontaneous magnetization $M(T)$, including what is referred to in this paper as “SW-MFA” minimal model for the single-sublattice ferromagnet. In contrast to YIG, this approach describes the experimental data in EuO and EuS quite well. The main reason for this success is the small value of the applicability parameter $1/Z\approx 1/12$.
Additionally, the simple crystallographic structures of EuO and EuS allow us to develop a suitable diagrammatic approach to calculate the $1/Z$ corrections to SW-MFA. As a result, we obtained excellent quantitative agreement between our theoretical predictions (which did not involve any fitting parameters) and experimental data, as noted in Ref. [6].
Unfortunately, at this stage, we cannot fully apply this method in the study of YIG. The absence of a small applicability parameter for the SW-MFA approach in YIG (where $Z\approx 5$) leads to only a qualitative agreement between the SW-MFA approach and experimental results. Additionally, the complexity of YIG’s 20-sublattice crystallographic structure makes the analytical calculation of $1/Z$ corrections difficult and even impractical, as $1/Z$ is not small. In view of these challenges, we limited our study of YIG at the SW-MFA level, achieving quantitative agreement with experiments by introducing a tuning parameter $\delta$.
Appendix C Positions of Fe3+ ions in elementary cell of YIG
To define the number of nearest neighbors and the relevant distances to assign the coordination spheres, we have drawn in Fig. 1 (b) the Fe3+ ions, oriented such that (a)-Fe3+ ions are placed in the corners and the center of the half-cube. The particular Fe3+ ion located at (a)-position (0,0,0) (see coordinate labels in Fig. 1 (a)) has $Z_{\rm aa}=8$ nearest-neighbor (a)-Fe3+ ions in positions $(\pm\frac{1}{4},\pm\frac{1}{4},\pm\frac{1}{4})$, separated by distance $\Delta_{\rm aa}=\dfrac{\sqrt{3}}{4}\approx 0.433$. The coordinates of the nearest $Z_{\rm ad}=6$ neighboring sites (d) with respect to the (a)-Fe3+ ion at $(0,0,0)$ are ($0,\frac{1}{4},-\frac{1}{8}$), ($0,-\frac{1}{4},\frac{1}{8}$); ($\frac{1}{8},0,-\frac{1}{4}$), ($\frac{1}{8},0,\frac{1}{4}$); ($\frac{1}{4}$, $\frac{1}{8}$, 0), ($\frac{1}{4},\frac{1}{8},0$), with the separation $\Delta_{\rm ad}=\dfrac{\sqrt{5}}{8}\approx 0.2795$.
[…]
Appendix D Exchange integrals
@@
6 Princep [28] 2017 Improved Neutron scattering of SW -79    0.0 $|$ -12.8 -6.0 1018 -0.82 86
@@
Table 1: Imporant parameters of YIG. The columns from left to right: the reference to the data origin and the method used to extract it; the exchange integrals $J_{ij}$ (values in K), the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF}}$ in the mean-field approximation computed using Eq. (16a); the ratio $\overline{S}_{\rm d}^{z}/\overline{S}_{\rm a}^{z}$ for $T\to T_{{}_{\rm C}}^{{}^{\rm MF}}$ from below, Eq. (16b); the exchange frequency $\omega_{\rm ex}$ (in K) according to (42c). Estimates of $\omega_{\rm ex}$ obtained by various authors, using a number of techniques, vary from 40 to 100 K, see Tab. 1 in [53]. The experimental value of YIG Curie temperature is $T_{{}_{\rm C}}^{\rm Exp}\approx 559\,$K. For the comparison with other published data, note that 11\,meV $\approx 11.6\,$K.
@@
Appendix E MFA with dipole-dipole interaction
To characterize the effect of dipole-dipole interaction on the magnetization of magnetically ordered dielectrics in the MFA, we introduce a characteristic temperature
$T_{\rm dd}\equiv\frac{\mu_{{}_{\rm B}}M_{*}}{k_{{}_{\rm B}}}=0.067\,\mbox{K}\ .$ (34)
Here, $\mu_{{}_{\rm B}}=9.274\times 10^{-21}$erg/Oe is the Bohr magneton and the Boltsman constant, denoted as $k_{{}_{\rm B}}=1.381\times 10^{-16}\,$erg/K. $M_{*}$ is the typical value of the magnetization, taken for the estimate in the right-hand side of (34) equal to $M_{*}=1000$ Oe.
For example, in the Weiss MFA, which accounts only for dipole-dipole interaction of classical spins [7, 6], the Curie temperature
$T_{{}_{\rm C}}^{{}^{\rm W}}=4\pi N_{z}\mu_{{}_{\rm B}}M(0)=\frac{4\pi N_{z}}{3}\frac{M(0)}{M_{*}}T_{\rm dd}\ .$ (35)
Here $M(T)$ denotes the magnetization. At $T=0$, the magnetization in EuO is $M(0)\approx 1920\,$Oe, while in YIG, it is about 200 Oe. The demagnetization factor $N_{z}\leq 1$ depends on the sample’s shape. For spherical samples, $N_{z}=1/3$. see e.g. [55].
In the presence of magnetic dipole-dipole interactions, the effective exchange magnetic field gains an additional contribution known as the demagnetization magnetic field:
$H_{\rm dem}=-4\pi N_{z}M(T)\ .$ (36)
The dipole-dipole interactions in single-sublattice ferromagnets, such as Eu0 and EuS, renormalize the effective magnetic field $H^{\rm eff}$ given by (4a) as follows
$H^{\rm eff}\Rightarrow\widetilde{H}^{\rm eff}=H^{\rm eff}-H_{\rm dem}\mu_{{}_{\rm B}}\ .$ (37)
Appearance of $\mu_{{}_{\rm B}}$ follows from the fact that $H^{\rm eff}$, traditionally called effective (magnetic) field, actually is the energy of spin in some magnetic field.
Taking into account Eq. (4a) along with (1) and $M(0)=\mu_{{}_{\rm B}}S$ one concludes
$\widetilde{H}^{\rm eff}=\overline{S}^{z}\widetilde{\cal J}_{0}\,,\quad\widetilde{\cal J}_{0}={\cal J}_{0}-4\pi N_{z}\mu^{2}_{{}_{\rm B}}\ .$ (38)
Next, using Eq. (6), one finds the equation for the correction to the Curie temperature caused by the dipole-dipole interactions:
$\begin{split}\Delta T=&\frac{4\pi N_{z}S(S+1)}{3}\mu^{2}_{{}_{B}}=\frac{4\pi N_{z}(S+1)}{3}M(0)\mu_{{}_{\rm B}}\\ =&\frac{4\pi N_{z}(S+1)}{3}\frac{M(0)}{M_{*}}T_{\rm dd}\ .\end{split}$ (39)
An accurate calculation of the correction $\Delta T$ in multi-sublattice ferrimagnets gives the result (39) in which $S$ is the total spin of the elementary cell (equal to 10 in YIG) with an additional numerical factor close to unity.
Taking for EuO $M(0)\approx 1920\,$Oe and $S=7/2$, $N_{z}=1/3$ (for qubic sample), we find
$\Delta T\approx 2\pi\,T_{\rm dd}\frac{M(0)}{M_{*}}\approx 0.79\,\mbox{K}\ .$ (40)
A similar estimate for YIG with $M(0)\approx 200\,$Oe and $S=10$ gives
$\Delta T\approx\frac{44\pi}{9}T_{\rm dd}\frac{M(0)}{M_{*}}\approx 0.21\,\mbox{K}\,$, (41)
which is negligible compared to $T_{{}_{\rm C}}\approx 560\,$K.
Analyzing the temperature behavior of $\overline{S}^{z}(T)$ with the $\widetilde{H}^{\rm eff}$, given by (37) we conclude that the relative contribution of the dipole-dipole interaction is about $T\Delta T/T_{{}_{\rm C}}^{2}\ll 1$. Therefore, except for the small temperature range around $T_{{}_{\rm C}}$, in which the critical behavior is observed (as considered e.g. in [26, 25]), we can safely neglect the contribution of the dipole-dipole interaction to the MF behavior of YIG across the entire temperature region.
Appendix F Approximations for magnon frequencies $\omega_{j}(k)$
@@
[…]
in agreement with the well known result [32]. We see that the contribution of the exchange interaction to the frequency of the ferromagnetic branch (42c) vanishes in the long spin-wave limit $k\to 0$. In this limit, one has to account for the contribution of the magnetic dipole-dipole interaction, which is independent of the magnitude of $k$, being dependent only on the direction of vector $\bm{k}$ via $\Theta_{\bm{k}}$:
$\omega(\bm{k})=g(H_{0}-4\pi MN_{z}+2\pi M\sin^{2}\Theta_{\bm{k}})+\omega_{\rm ex}(ak)^{2}\,$, (43)
see, e.g. Eq.(3.1.10) in [55]. Here, $g$ is the gyromagnetic factor, $H_{0}$ is the value of the external magnetic field, $\Theta_{\bm{k}}$ is the angle between $\bm{k}$ and $\bm{M}$, and $N_{z}$, as before, is the demagnetization factor ($N_{z}=1/3$ for spherical samples).
Equation  (43) allows one to estimate the crossover wavenumber $k_{*}$ at which the contributions of the dipole-dipole and exchange interactions to the frequency become equal: $\omega_{\rm ex}(ak_{*})^{2}=4\pi gM$. In YIG at room temperature, this gives $ak_{*}\approx 0.07$[55]. This value is significantly smaller than the size of the Brillouin zone (with $ak\sim 5$, see Fig. 3). Consequently, the contribution of very long spin waves, where dipole-dipole interactions become important, can be safely ignored. This enables us to omit them from the integrals in (23) for $\cal N$$_{\pm}$, which determine, according to Eqs.(22), the spin-wave suppression of magnetization in our approach.
Returning back to (42c), notice that for $q>q_{*}$, above some relatively small transition value $q_{*}=1/\sqrt{25.4}\approx 0.2$, $\omega(k)$ becomes almost linear in $k$
$\omega_{\rm da1}\approx(-\omega_{\rm neg}+\omega_{\rm lin}\,ak)\,$, (44)
as seen in Fig. 3. In (44), $\omega_{\rm neg}\approx 110\,$K and $\omega_{\rm lin}\approx 170\,$K. The nearly linear part of the dispersion laws $\omega_{\rm da1}$ and $\omega_{\rm ad1}$ originates from the fact that YIG is close to the antiferromagnetic arrangement  [^2] In YIG, the ratio of the numbers of oppositely oriented spins is $12/8=3/2$. Their difference 128=412-8=4 is small with respect to half of their sum =10=10 ($4/10=0.2$). In this sense, the YIG magnetic structure is relatively close to AFM, where the dispersion laws exhibit a similar form [55].
@@
Appendix G Measuring Procedure and Apparatus
The magnetization measurements for this work were carried out using a Physical Property Measurement System (PPMS, Quantum Design), which allows temperature control from 400 K down to 1.8 K and enables the application of external magnetic fields of up to 90 kOe, generated by a pair of superconducting coils in Helmholtz configuration. The measurements were performed using the vibrating sample magnetometer (VSM) option integrated into the PPMS as an instrument insert.
A commercially obtained YIG sphere (Microspheres,Inc) with a diameter of 1 mm was fixed to the sample holder using nonmagnetic adhesive, with the easy axis—the $\langle 111\rangle$ crystallographic direction—aligned parallel to the applied external magnetic field. During the measurements, a constant external field of 10 kOe was applied, which is sufficient to ensure magnetic saturation of the YIG sphere across the entire available temperature range.
@@ Appendix G Measuring Procedure and Apparatus
Appendix H Numerical Methods
@@
For the fit shown in Fig. 2, represented by the violet curve ④, we used the experimental data from Ref. [21] [orange dots ②]. The high-temperature fit was based on the phenomenological expression $\overline{S}^{z}=A(T_{{}_{\rm C}}-T)^{\beta}$, where $A$ and $\beta$ are fit parameters. The fitting was performed using the NonlinearModelFit function in Mathematica. To ensure a reliable result, it was necessary to carefully select the temperature range for the fit. Data in the interval from 350 K to 535 K was chosen. The upper bound had to remain below the Curie temperature $T_{{}_{\rm C}}=560\,\text{K}$, as data too close to $T_{{}_{\rm C}}$ is affected by an experimental artifact—residual magnetization above $T_{{}_{\rm C}}$ caused by measurements taken in a relatively strong external magnetic field of 10 kOe. Conversely, the lower bound was set near room temperature, as the fitting model does not account for the saturation of magnetization at lower temperatures; thus, including such data would distort the fit parameters. It is worth noting that the fit results remained consistent under variations of the chosen interval limits by $\pm 10\,\text{K}$.
The curves $\overline{S}^{z}_{\mathrm{a}}(T)/\overline{S}^{z}_{\mathrm{a}}(0)]$ ⑨ and $\overline{S}^{z}_{\mathrm{d}}(T)/\overline{S}^{z}_{\mathrm{d}}(0)$ ⑩ in Fig. 4, as well as their difference $\overline{S}^{z}(T)/\overline{S}^{z}(0)$ ⑤ in Figs. 2 and 4, were obtained by numerically solving the system of Eqs. (15) using the parameters specified in Eq. (14). The numerical solution was carried out using the FindRoot function in Mathematica.
For the curves ⑦ in Fig. 2 and the additional curves ⑪ and ⑫ in Fig. 5, Eq. (22a) was evaluated. This required calculating the total number of magnons within the first Brillouin zone as given by Eq. (22b), which involves an integration over the wavevector $\bm{k}$. In this work, we approximate the first Brillouin zone by a sphere of radius equal to the Debye wavevector $k_{\rm D}$, which for a BCC lattice is given by $k_{\rm D}=\sqrt[3]{12\pi^{2}/a^{3}}$, a simplification that provides a good approximation for our purposes. Accordingly, the integral defined in Eq. (23) over $n_{j}(\bm{k})$, Eq. (24), was performed in spherical coordinates, using the magnon dispersion relations $\omega_{j}(\bm{k})$ provided in Appendix F. The integration range was set from $k_{\rm min}=0$ to $k_{\rm max}=k_{\rm D}$. The numerical evaluation was carried out using the NIntegrate function in Mathematica.